The event of you going to work is A and the event of you taking leave is B. If these events are mutually exclusive events, using P(A)=0.55, and P(B)=0.10, what is P(A|B)?
Answer #1

P(A/B)

= P(A and B) / P(B) (Using conditional probability rule)

It is given that A and B are mutually exclusive events

Thus, P(A and B) = 0

Thus, p(A/B) = 0/p(B)

= 0